# Dyn-HTE: High-temperature expansion of the dynamic Matsubara spin correlator

### v1

@@ IV.2 Graph-based evaluation of bond-sums
Compared to the established HTE for equal-time spin-spin correlators $\left\langle S_{i}^{z}S_{i^{\prime}}^{z}\right\rangle$, the new and challenging aspect in Eq. (IV.1) are the $n+2$-dimensional imaginary time integrals. Relegating the integral evaluation to the end of this section we first focus on the bond-sums exposed in Eq. (14) for which efficient evaluation strategies are well developed [2]. The following discussion of the necessary elements of these strategies is self-contained and does not assume any prior knowledge. In particular we adopt the graph-based approach and re-organize the sum over lattice bonds $\sum_{b_{1},...,b_{n}}$ as a sum over evaluations of $F^{(n)}_{ii^{\prime}}$ on abstract (multi-)graphs $g^{(n)}$ which we will define momentarily,
@@
A (multi-)graph of order $n$ denoted by $g^{(n)}$ is defined by the multi-set of $n$ (not necessarily distinct) edges $\{e_{1},...,e_{n}\}$ connecting its (arbitrarily numbered) vertices which contain the terminal vertices $j,j^{\prime}$. Different graphs are distinguished by graph topology, see Fig. 1 for the example of all distinct graphs $g^{(n)}_{t}$, $t=1,2,...$ with up to $n=3$ edges that are required for our purpose. Vertices are indicated by blue circles, the terminal vertices $j,j^{\prime}$ are connected to red-square terminal indicators by thin lines. As we motivate later, two graphs only differing by the exchange $j\leftrightarrow j^{\prime}$ are considered topologically equivalent.
Figure 1: All required graphs $g^{(n)}_{t}$ with $n=0,1,2,3$ edges (black lines) and arbitrary numbered vertices (blue circles). Terminal vertices $j,j^{\prime}$ are indicated by their red-square terminal flags attached with thin gray lines. The symmetry factor is $s[g^{(n)}_{t}]=1$ for all graphs shown except for $(n,t)=(3,5)$ where it is two (exchange of vertices $2\leftrightarrow 3$).
The next concept in Eq. (15) is the embedding factor $e(\mathcal{L},i,i^{\prime},g^{(n)})\in\{0,1,2,...\}$. It counts the number of sub-graph isomorphisms from the graph $g^{(n)}$ (with edge-multiplicities ignored) to the lattice $\mathcal{L}$ divided by the graphs’ symmetry factor $s[g^{(n)}]\in\{1,2,...\}$ to avoid overcounting due to the unphysical vertex numbering. The symmetry factor is the number of graph isomorphisms that keep the terminal vertices invariant and respect the edge-multiplicities, see the caption of Fig. 1 for examples for $s[g^{(n)}]$.
For the embedding $g^{(n)}\rightarrow\mathcal{L}$ we require $(i,i^{\prime})\rightarrow(j,j^{\prime})$ or $(i,i^{\prime})\rightarrow(j^{\prime},j)$. Here, the latter possibility, which allows us to avoid graphs connected to another one by an exchange of $j\leftrightarrow j^{\prime}$, is not counted if $g^{(n)}$ is mapped to itself under this operation. As an example, embedding factors for a number of simple lattices are given in Tab. 1. For two- and in particular three-dimensional lattices the embedding step can be numerically demanding. To ensure efficiency, our numerical implementation [21] builds on lattice generating functionalities provided in the Monte-Carlo software package SpinMC.jl [49] and also relies on space-group symmetries for speed-up.
@@
Table 1: Embedding factor $e(\mathcal{L},i,i^{\prime},g^{(n)}_{t})$ for various lattices $\mathcal{L}$ and various separations between external sites $i,i^{\prime}$. Here ’loc’ means local ($i=i^{\prime}$) while ’(n)nn’ stands for (next-)nearest neighbor. A missing entry means $e(\mathcal{L},i,i^{\prime},g^{(n)}_{t})=0$.
@@
Following these considerations, we have created all required graphs $g^{(n)}$ [with a potential non-zero graph-evaluation $c_{g^{(n)}}(i\nu_{m})$], for $n=0,1,...,12$. The number of graphs per order $n\gtrsim 6$ is roughly a factor $4$ larger than for the previous order. At our highest order $n=12$ we need to consider 1273854 graphs $g^{(12)}$.
The final quantity to be discussed in Eq. (15) are the graph evaluations $c_{g^{(n)}}$ of Eq. (16). The sum $\sum_{\{e_{1},...,e_{n}\}\rightarrow\{\tilde{e}_{1},...,\tilde{e}_{n}\}}$ associates the multi-set of the graph’s edges to the time-ordered bond operators in $F^{(n)}_{jj^{\prime}}\left(\tilde{e}_{1},...,\tilde{e}_{n};i\nu_{m}\right)$ in all $n!/d[g^{(n)}]$ different ways. Here $d[g^{(n)}]$ yields the graph degeneracy which is the product of factorials of all edge weights.
@@
where $\Delta_{y}\equiv(1-\delta_{0,y})/y$. As we argue later, after the embedding, all odd $l$ vanish on the right-hand side of Eq. (15) and are thus not computed in the first place. The $c^{(n)}_{t;l}$ for $l=0,2,...,r_{max}$ are rational numbers and together with Eq. (10) they give rise to the final form (7). In Tab. 2 we provide the values for the evaluations $c^{(n)}_{t;l}$ for all graphs $g^{(n)}$ with $n\leq 3$ shown in Fig. 1. The remainder of this section is concerned with the actual calculation of these graph evaluations.
@@ IV.3 Recursive sub-graph subtractions
For the evaluation of the dynamic spin correlator on graph $g^{(n)}$ denoted by $c_{g^{(n)}}(i\nu_{m})$ and defined in Eq. (16) we are required to assign the multi-set of the graph’s edges $\{e_{1},...,e_{n}\}$ to the bond-indices of the set of operators $\{V_{\tilde{e}_{1}}(\tau_{1}),...,V_{\tilde{e}_{n}}(\tau_{n})\}$ according to $\{e_{1},...,e_{n}\}\rightarrow\{\tilde{e}_{1},...,\tilde{e}_{n}\}$ in all possible ways. The operators are distinguishable by their time arguments even for identical edges. From Eq. (IV.1), we obtain two terms:
@@
$c_{g^{(n)}}^{[full]}(i\nu_{m})=\frac{T^{n+1}}{d[g^{(n)}]}\int_{0}^{\beta}e^{i\nu_{m}(\tau-\tau^{\prime})}\mathrm{d}\tau_{1}...\mathrm{d}\tau_{n}\mathrm{d}\tau\mathrm{d}\tau^{\prime}$
×$\times\left\langle\mathcal{T}V_{e_{1}}(\tau_{1})...V_{e_{n}}(\tau_{n})S_{j}^{z}(\tau)S_{j^{\prime}}^{z}(\tau^{\prime})\right\rangle_{0},$ (19)
@@
For the second “subtracted” term in Eq. (18), we rewrite the $\sum_{S\subsetneq\{1,...,n\}}$ in Eq. (IV.1) as a sum over subgraphs $g^{(k)}\subsetneq g^{(n)}$ with $k<n$ edges. The subgraph $g^{(k)}=\{e_{1}^{\prime},...,e_{k}^{\prime}\}$ still needs to be connected and must contain the terminal vertices $j,j^{\prime}$. This leaves the vacuum graph without terminals $g^{(n)}\backslash g^{(k)}$ which is possibly disconnected. There can be more than one ways to get to the same $g^{(k)}$, this multiplicity factor is denoted by $f(g^{(n)},g^{(k)})\in\{1,2,...\}$. We refer to Fig. 2 for an example showing all possible non-zero subtractions from a particular graph $g^{(6)}$ with $n=6$ edges (left) shown in the 2nd and 3rd column together with their factors $f(g^{(n)},g^{(k)})$ and the $g^{(k)}$ (top), the vacuum graph is shown in the bottom. Finally, in the the second line of Eq. (IV.1) there are $\binom{n}{k}$ possibilities to split the imaginary time labels between the first and second average. Hence we obtain
@@
$=\frac{\binom{n}{k}}{n!}\!\!\sum_{g^{(k)}\subsetneq g^{(n)}}f(g^{(n)},g^{(k)})T^{n+1}\!\!\int_{0}^{\beta}\!\!e^{i\nu_{m}(\tau-\tau^{\prime})}\mathrm{d}\tau_{1}...\mathrm{d}\tau_{n}\mathrm{d}\tau\mathrm{d}\tau^{\prime}$
×$\times\!\!\!\!\sum_{g^{(k)}=\{e_{1}^{\prime},...,e_{k}^{\prime}\}\rightarrow\{\tilde{e}^{\prime}_{1},...,\tilde{e}^{\prime}_{k}\}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!\left\langle\mathcal{T}V_{\tilde{e}^{\prime}_{1}}(\tau_{1})...V_{\tilde{e}^{\prime}_{k}}(\tau_{k})S_{j}^{z}(\tau)S_{j^{\prime}}^{z}(\tau^{\prime})\right\rangle_{0}^{\!\mathrm{V-con}}$
×$\times\!\!\!\!\sum_{g^{(n)}\backslash g^{(k)}=\{e^{\prime}_{k+1},...,e^{\prime}_{n}\}\rightarrow\{\tilde{e}^{\prime}_{k+1},...,\tilde{e}^{\prime}_{n}\}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!\left\langle\mathcal{T}V_{\tilde{e}^{\prime}_{k+1}}(\tau_{k+1})...V_{\tilde{e}^{\prime}_{n}}(\tau_{n})\right\rangle_{0}.$
We obtain a recursive equation for $c_{g^{(n)}}(i\nu_{m})$ by inserting the graph-resolved version of Eq. (11) for $c_{g^{(k)}}(i\nu_{m})$ into the right-hand side of Eq. (20). This yields
$c_{g^{(n)}}^{[sub]}(i\nu_{m})\!=\!\!\sum_{g^{(k)}\subsetneq g^{(n)}}f(g^{(n)},g^{(k)})\,c_{g^{(k)}}(i\nu_{m})\,\left\langle g^{(n)}\backslash g^{(k)}\right\rangle_{0},$ (21)
where the last term is the evaluation of the vacuum graph $g^{(n)}\backslash g^{(k)}$ defined as
$\left\langle g^{(n)}\backslash g^{(k)}\right\rangle_{0}=\frac{T^{n-k}}{(n-k)!}\int_{0}^{\beta}\mathrm{d}\tau_{k+1}...\mathrm{d}\tau_{n}$ (22)
×$\times\!\!\!\!\!\!\sum_{g^{(n)}\backslash g^{(k)}=\{e^{\prime}_{k+1},...,e^{\prime}_{n}\}\rightarrow\{\tilde{e}^{\prime}_{k+1},...,\tilde{e}^{\prime}_{n}\}}\!\!\!\!\!\!\left\langle\mathcal{T}V_{\tilde{e}^{\prime}_{k+1}}(\tau_{k+1})...V_{\tilde{e}^{\prime}_{n}}(\tau_{n})\right\rangle_{0}$
@@
In the last step, time-integrals together with the time-ordering $\mathcal{T}$ lead to the sum over all permutations $p$ of the numbers $1,2,...,n-k$, similar to vacuum graph evaluation for conventional HTE. For disconnected vacuum graphs $g^{(n)}\backslash g^{(k)}\equiv v$ which separate into two connected components $v=v_{1}\dot{\cup}v_{2}$ we have $\left\langle v\right\rangle_{0}=\left\langle v_{1}\dot{\cup}v_{2}\right\rangle_{0}=\left\langle v_{1}\right\rangle_{0}\left\langle v_{2}\right\rangle_{0}$ and this generalizes to an arbitrary number of connected components.
In summary, once all the “full” graph evaluations (19) and vacuum graph evaluations (23) of orders smaller or equal to $n-1$ have been calculated, Eqns. (18) and (21) can be used to find the $n$-th order graph evaluations via the recursion
@@
Crucially, as we are here dealing with spin operators, Wick’s theorem is not applicable and the first term on the right-hand side of Eq. (IV.3), $c_{g^{(n)}}^{[full]}(i\nu_{m})$ given in Eq. (19), cannot be written as a determinant or permanent. On the other hand, it turns out that the time-dependence of the integrand in our case is much simpler than that encountered in the Hubbard model. These observations lead to a dedicated strategy which we discuss in the next subsection.

### v2

@@ IV.2 Graph-based evaluation of bond-sums
Compared to the established HTE for equal-time spin-spin correlators $\left\langle S_{i}^{z}S_{i^{\prime}}^{z}\right\rangle$, the new and challenging aspect in Eq. (IV.1) are the $n+2$-dimensional imaginary time integrals. We relegate the integral evaluation to section IV.4. Here we first focus on the bond-sums in Eq. (IV.1), which we expose by summarizing the rest as follows,
$c_{ii^{\prime}}^{(n)}(i\nu_{m})\equiv\sum_{b_{1},...,b_{n}}F^{(n)}_{ii^{\prime}}\left(b_{1},...,b_{n};i\nu_{m}\right).$ (14)
Efficient evaluation strategies for these bond-sums are well developed in the literature on conventional HTE [2]. The following discussion of the necessary elements of these strategies is self-contained and does not assume any prior knowledge. In a nutshell, our plan is to evaluate the expansion of the Matsubara correlator not directly for a full lattice $\mathcal{L}$, but first for individual lattice snippets with only $n$ bonds which are called graphs, $g^{(n)}$. These are then added up as they fit into the full lattice. Formally speaking, we re-organize the sum over (potentially identical) lattice bonds $\sum_{b_{1},...,b_{n}}$ as a sum over evaluations of $F^{(n)}_{ii^{\prime}}$ on these graphs,
@@
A (multi-)graph of order $n$ denoted by $g^{(n)}$ can be thought of as a lattice-snippet with $n$ bonds, see Fig. 1. Formally, it is defined by the multi-set of $n$ (not necessarily distinct) edges $\{e_{1},...,e_{n}\}$ (thick black lines) connecting its arbitrarily numbered vertices (blue circles) which contain the terminal vertices $j,j^{\prime}$ (blue circles highlighted by attached red-square flags). The terminal vertices indicate the position of the external operators $S_{i}^{z}$ and $S_{i^{\prime}}^{z}$. Different graphs are distinguished by graph topology, Fig. 1 shows all topological distinct graphs $g^{(n)}_{t}$, labeled by $t=1,2,$... with up to $n=3$ edges that are required for our purpose. As we motivate later, two graphs only differing by the exchange of terminals $j\leftrightarrow j^{\prime}$ are considered topologically equivalent.
Figure 1: All required graphs $g^{(n)}_{t}$ with $n=0,1,2,3$ edges (black lines) and arbitrary numbered vertices (blue circles). Terminal vertices $j,j^{\prime}$ (also blue circles) are indicated by their red-square terminal flags attached with thin gray lines. The symmetry factor is $s[g^{(n)}_{t}]=1$ for all graphs shown except for $(n,t)=(3,5)$ where it is two (exchange of vertices $2\leftrightarrow 3$).
Having defined the graph $g^{(n)}$ as a lattice snippet, we next need to place it into the full lattice $\mathcal{L}$. This is done at the embedding step, formally expressed in Eq. (15). Here the embedding factor $e(\mathcal{L},i,i^{\prime},g^{(n)})\in\{0,1,2,$...\} counts the number of sub-graph isomorphisms from the graph $g^{(n)}$ (with edge-multiplicities ignored) to the lattice $\mathcal{L}$ divided by the graphs’ symmetry factor $s[g^{(n)}]\in\{1,2,$...\} to avoid overcounting due to the arbitrary vertex numbering. The symmetry factor is the number of graph isomorphisms that keep the terminal vertices invariant and respect the edge-multiplicities, see the caption of Fig. 1 for examples for $s[g^{(n)}]$.
For the embedding $g^{(n)}\rightarrow\mathcal{L}$ we require the assignments $(i,i^{\prime})\rightarrow(j,j^{\prime})$ or $(i,i^{\prime})\rightarrow(j^{\prime},j)$ to match the positions of the external operators. Taking into account also the latter assignment $(i,i^{\prime})\rightarrow(j^{\prime},j)$ allows us to skip graphs which are connected to another graph in the list by an exchange of $j\leftrightarrow j^{\prime}$. This assignment is not counted if $g^{(n)}$ is mapped to itself under $j\leftrightarrow j^{\prime}$. As an example, embedding factors for a number of simple lattices are given in Tab. 1. For two- and in particular three-dimensional lattices the embedding step can be numerically demanding. To ensure efficiency, our numerical implementation [21] builds on lattice generating functionalities provided in the software package SpinMC.jl [47], relies on space-group symmetries for speed-up and uses advanced graph-theoretical algorithms [48, 49].
@@
Table 1: Embedding factor $e(\mathcal{L},i,i^{\prime},g^{(n)}_{t})$ for various lattices $\mathcal{L}$ and various separations between external sites $i,i^{\prime}$. Here ’loc’ means local ($i=i^{\prime}$) while ’(n)$n$’ stands for (next-)nearest neighbor. A missing entry means $e(\mathcal{L},i,i^{\prime},g^{(n)}_{t})=0$.
@@
Following these considerations, we have created lists of all required graphs $g^{(n)}$ [with a potential non-zero graph-evaluation $c_{g^{(n)}}(i\nu_{m})$], for n=0,1,,12n=0,1,...,12. The number of graphs per order $n\gtrsim 6$ is roughly a factor $4$ larger than for the previous order. At our highest order $n=n_{max}=12$ we need to consider 1273854 graphs $g^{(12)}$.
Finally we need to consider the quantity which we wanted to evaluate on the lattice snippets defined by a selection of edges $\{e_{1},...,e_{n}\}$: The HTE of the Matsubara correlator. We call this HTE the graph evaluation $c_{g^{(n)}}(i\nu_{m})$. In Eq. (15) the correlator expansion on the full lattice is obtained by a sum over all graphs evaluations weighted with the embedding factors. In Eq. (16), the individual graph evaluations are given by a sum $\sum_{\{e_{1},...,e_{n}\}\rightarrow\{\tilde{e}_{1},...,\tilde{e}_{n}\}}$ which associates the multi-set of the graph’s edges to the time-ordered bond operators in $F^{(n)}_{jj^{\prime}}\left(\tilde{e}_{1},...,\tilde{e}_{n};i\nu_{m}\right)$ in all $n!/d[g^{(n)}]$ different ways. Here $d[g^{(n)}]$ yields the graph degeneracy which is the product of factorials of all edge weights.
@@
where $\Delta_{2\pi m}\equiv(1-\delta_{0,m})/(2\pi m)$ is the inverse of dimensionless (non-zero) Matsubara frequencies. As we argue later, after the embedding, all odd $l$ vanish on the right-hand side of Eq. (15) and are thus not computed in the first place. The $c^{(n)}_{t;l}$ for $l=0,2,...,r_{max}$ are rational numbers and together with Eq. (10) they give rise to the final form (7). In Tab. 2 we provide the values for the evaluations $c^{(n)}_{t;l}$ for all graphs $g^{(n)}$ with $n\leq 3$ shown in Fig. 1. The remainder of this section is concerned with the actual calculation of these graph evaluations.
@@ IV.3 Recursive sub-graph subtractions
For the (graph-)evaluation of the dynamic spin correlator on graph $g^{(n)}$ denoted by $c_{g^{(n)}}(i\nu_{m})$ in Eq. (16) we are required to assign the multi-set of the graph’s edges $\{e_{1},...,e_{n}\}$ to the bond-indices of the operators $\{V_{\tilde{e}_{1}}(\tau_{1}),...,V_{\tilde{e}_{n}}(\tau_{n})\}$ according to $\{e_{1},...,e_{n}\}\rightarrow\{\tilde{e}_{1},...,\tilde{e}_{n}\}$ in all possible ways. The operators are distinguishable by their time arguments even for identical edges. However, recall the iterative definition of the $\mathrm{V-con}$ correlator in Eq. (IV.1) which carries over to the graph evaluation as follows:
@@
$c_{g^{(n)}}^{[full]}(i\nu_{m})=\frac{T^{n+2}}{d[g^{(n)}]}\int_{0}^{\beta}e^{i\nu_{m}(\tau-\tau^{\prime})}\mathrm{d}\tau_{1}...\mathrm{d}\tau_{n}\mathrm{d}\tau\mathrm{d}\tau^{\prime}$
$\times\left\langle\mathcal{T}V_{e_{1}}(\tau_{1})...V_{e_{n}}(\tau_{n})S_{j}^{z}(\tau)S_{j^{\prime}}^{z}(\tau^{\prime})\right\rangle_{0},$ (19)
@@
For the second “subtracted” term in Eq. (18), we rewrite the $\sum_{S\subsetneq\{1,...,n\}}$ in Eq. (IV.1) as a sum over subgraphs $g^{(k)}\subsetneq g^{(n)}$ with $k<n$ edges. The subgraph $g^{(k)}=\{e_{1}^{\prime},...,e_{k}^{\prime}\}$ still needs to be connected and must contain the terminal vertices $j,j^{\prime}$. This remaining edges $g^{(n)}\backslash g^{(k)}$ form a vacuum graph without external vertices which is possibly disconnected. There can be more than one ways to get to topologically equivalent $g^{(k)}$, this multiplicity factor is denoted by $f(g^{(n)},g^{(k)})\in\{1,2,$...\}. We refer to Fig. 2 for an example listing all possible non-zero subtractions from a particular graph $g^{(6)}$ with $n=6$ edges (left) shown in the 2nd and 3rd column together with their factors $f(g^{(n)},g^{(k)})$ and the $g^{(k)}$ (top), the vacuum graph $g^{(n)}\backslash g^{(k)}$ is shown in the bottom. Finally, in the second line of Eq. (IV.1) there are $\binom{n}{k}$ possibilities to distribute the imaginary time labels between the first and second average. Hence we obtain
@@
$=\frac{\binom{n}{k}}{n!}\!\!\sum_{g^{(k)}\subsetneq g^{(n)}}\!\!\!f(g^{(n)},g^{(k)})T^{n+2}\!\!\int_{0}^{\beta}\!\!e^{i\nu_{m}(\tau-\tau^{\prime})}\mathrm{d}\tau_{1}...\mathrm{d}\tau_{n}\mathrm{d}\tau\mathrm{d}\tau^{\prime}$
$\times\!\!\!\!\sum_{g^{(k)}=\{e_{1}^{\prime},...,e_{k}^{\prime}\}\rightarrow\{\tilde{e}^{\prime}_{1},...,\tilde{e}^{\prime}_{k}\}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!\left\langle\mathcal{T}V_{\tilde{e}^{\prime}_{1}}(\tau_{1})...V_{\tilde{e}^{\prime}_{k}}(\tau_{k})S_{j}^{z}(\tau)S_{j^{\prime}}^{z}(\tau^{\prime})\right\rangle_{0}^{\!\mathrm{V-con}}$
$\times\!\!\!\!\sum_{g^{(n)}\backslash g^{(k)}=\{e^{\prime}_{k+1},...,e^{\prime}_{n}\}\rightarrow\{\tilde{e}^{\prime}_{k+1},...,\tilde{e}^{\prime}_{n}\}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!\left\langle\mathcal{T}V_{\tilde{e}^{\prime}_{k+1}}(\tau_{k+1})...V_{\tilde{e}^{\prime}_{n}}(\tau_{n})\right\rangle_{0}.$
The sum in the penultimate line assigns the multi-set of graph $g^{(k)}$’s edges $\{e_{1}^{\prime},...,e_{k}^{\prime}\}$ in all possible ways to the bond-indices of the operators $V_{\tilde{e}^{\prime}_{1}}(\tau_{1})$$V_{\tilde{e}^{\prime}_{k}}(\tau_{k})$ and similar in the last line. We now resolve these assignment-sums.
The penultimate line in Eq.(20) appears in the graph-resolved version of Eq. (11) (which only sums over the edges of graph $g^{(k)}$ instead of all bonds $b_{1},...,b_{n}$ of the lattice). Hence we obtain a recursive equation for the graph-evaluation $c_{g^{(n)}}(i\nu_{m})$
$c_{g^{(n)}}^{[sub]}(i\nu_{m})\!=\!\!\sum_{g^{(k)}\subsetneq g^{(n)}}f(g^{(n)},g^{(k)})\,c_{g^{(k)}}(i\nu_{m})\,\left\langle g^{(n)}\backslash g^{(k)}\right\rangle_{0}.$ (21)
The last term is the evaluation of the vacuum graph $g^{(n)}\backslash g^{(k)}$ defined from the remaining terms as
$\left\langle g^{(n)}\backslash g^{(k)}\right\rangle_{0}=\frac{T^{n-k}}{(n-k)!}\int_{0}^{\beta}\mathrm{d}\tau_{k+1}...\mathrm{d}\tau_{n}$
$\times\!\!\!\!\!\!\sum_{g^{(n)}\backslash g^{(k)}=\{e^{\prime}_{k+1},...,e^{\prime}_{n}\}\rightarrow\{\tilde{e}^{\prime}_{k+1},...,\tilde{e}^{\prime}_{n}\}}\!\!\!\!\!\!\left\langle\mathcal{T}V_{\tilde{e}^{\prime}_{k+1}}(\tau_{k+1})...V_{\tilde{e}^{\prime}_{n}}(\tau_{n})\right\rangle_{0}$
@@
In the last step, time-integrals together with the time-ordering $\mathcal{T}$ lead to the sum over all permutations $p$ of the numbers 1,2,,nk1,2,...,n-k, similar to vacuum graph evaluation for conventional HTE. For disconnected vacuum graphs $g^{(n)}\backslash g^{(k)}\equiv v$ which separate into two connected components $v=v_{1}\dot{\cup}v_{2}$ we have $\left\langle v\right\rangle_{0}=\left\langle v_{1}\dot{\cup}v_{2}\right\rangle_{0}=\left\langle v_{1}\right\rangle_{0}\left\langle v_{2}\right\rangle_{0}$ and this generalizes to an arbitrary number of connected components.
In summary, the recursive formulation necessitated by the removal of $V$-disconnected bond operator configurations in the definition (IV.1) works as follows: Once all the “full” graph evaluations (19) at order $n$ and vacuum graph evaluations (22) of orders smaller or equal to n1n-1 have been calculated, Eqns. (18) and (21) can be used to find the $n$-th order graph evaluations via the recursion
@@
Crucially, as we are here dealing with spin operators, Wick’s theorem is not applicable and the first term on the right-hand side of Eq. (IV.3), $c_{g^{(n)}}^{[full]}(i\nu_{m})$, given in Eq. (19), cannot be written as a determinant or permanent. On the other hand, it turns out that the time-dependence of the integrand in our case is much simpler than that encountered in the Hubbard model. These observations lead to a dedicated evaluation strategy which we discuss in the next subsection.